Mathematics 242 – Principles of Analysis Definitions Quiz 5 – March 14, 2014

Your Name:

B) (5) Define: a subsequence of a given sequence {x_{n}}.

A) (5) Give an example of a sequence that has three convergent subsequences, each converging to a different limit.